OpenStudy (anonymous):

Together a skateboard and a bicycle cost $199.00, the skateboard cost $51.00 less than the bicycle. How much does each cost?

OpenStudy (anonymous):

Pick a name for each of your variables.

OpenStudy (anonymous):

skateboard-74$ & bicycle-a25$

OpenStudy (anonymous):

Not helpful

OpenStudy (anonymous):

two equations will be x+y=199 & y=x-51 x is cost of bicycle & y is cost of skateboard

OpenStudy (anonymous):

more helpful

OpenStudy (anonymous):

solve it an get the answer

OpenStudy (anonymous):

*bicycle 125$

OpenStudy (anonymous):

let s=skateboard let b= bicycle s+b=199 s=b-51 b-51+b=199 2b=199+51 2b=250 b=125 s=125-51=74 74+125=199 so bicycle is $125 and skateboard is $74